# A Fictional Q&A Dataset for Studying Memorization and Knowledge Acquisition

### v1

@@ Abstract
When language models are trained on textual data, they acquire both knowledge about the structure of language as well as knowledge of facts about the world. At inference time, their knowledge of facts can be leveraged to solve interesting problems and perform useful knowledge work for users. It is well known that language models can verbatim memorize long sequences from their training data. However, it is much less well understood how language models memorize facts seen during training. In this work, we propose a new dataset to specifically empower researchers to study the dual processes of fact memorization and verbatim sequence memorization. The dataset consists of synthetically-generated, webtext-like documents about fictional events, as well as question-answer pairs about the events. We conduct training experiments showing how synthetic data about fictional events can be effective in teasing apart different forms of memorization. We also document the challenges in effectively building realistic, fictional synthetic data.
Figure 1: An illustration of the hierarchical structure of our fictional dataset. Small liberties taken in cropping and whitespace of the example texts for visualization purposes.
@@ 1 Introduction
It is well-known that language models memorize some of their training data. Sometimes memorization takes the form of verbatim memorization where exact sequences of tokens seen during training are likely to be outputted by the large language model (LLM). Vertbatim memorization ranges from the memorization of short common phrases (e.g. “the cat’s out of the bag”) to multi-paragraph excerpts from books or articles. Factual memorization is another form of memorization, in which facts about the world (e.g that cats see better in the dark than humans because their eyes have more rods) are learned as representations that can generalize to diverse downstream tasks. While sequence memorization may or may not be desirable depending on the length and nature of the sequence the LLM has memorized, generalizable fact memorization is almost always considered a desirable trait in LLMs. [^1] This poses challenges for trying to apply unlearning techniques to remove individual atoms of knowledge. Additionally, when models posses such capabilities, an inherent risk is copyright infringement. That being said, verbatim memorization, or exact reconstruction of training data is the primary issue for legal and copyright risks, not fact memorization. As our focus in this work is the latter, we do not discuss these topics any further in this work. See Lee et al. (2023); Cooper and Grimmelmann (2024) for a nuanced treatment of generative models and intellectual property. For example, user might reasonably expect to be able to ask an LLM “Why can cats see so well in the dark?” and get a correct answer, even if the knowledge to answer this question was only ever seen during training as part of a Wikipedia-style article about cat eyes.
@@
The phenomenon of verbatim memorization has been well studied; the work by  Carlini et al. (2019) serving as a canonical example in the domain of language models. However, we understand less about how language models memorize facts such that they are capable of using a learned fact for novel tasks at inference time. One challenge with studying the process of fact memorization during training is that it is very difficult to quantify how often a fact actually occurs during training. Prior work has studied the correlation between how well an LLM can answer questions about named entities with the frequency the named entity occurs in the training data (Kandpal et al., 2023). Others have trained very small models exclusively on synthetic biographies and then measured when the ability to answer biographical questions appears during model training (Allen-Zhu and Li, 2023a). Prior work has also sought to insert canaries (e.g. social security numbers or email addresses) during training and then check whether the model is capable of generating the canary string (Carlini et al., 2022b). No prior work has studied the training dynamics of LLMs acquiring generalizable knowledge of facts (i.e. fact memorization) under realistic large language models training settings. In this paper, we demonstrate how realistic, synthetic data about fictional events can be used to study the training dynamics of both fact and sequence memorization.
@@
Our primary contribution is the development of a dataset generation pipeline for producing corpora of documents about realistic but fictional events. While the textual styles and the statistical distribution of words and phrases in our data are similar to that of a natural pretraining corpus, we construct prompts which produce events with made-up people, places, and events. These dual characteristics of realism at the surface-level and fantasy at the content-level enable us to study the traits leading to memorization in a laboratory setting, with greater assurance that the facts contained in the data do not interact with any other knowledge in the pretraining corpus. In addition, the data pipeline we propose is unique in that it is a “living asset,” meaning that we can regenerate a fresh dataset for future experiments, and other researchers can tweak and repurpose parts of the recipe to suit their needs and explore other research questions than the ones we specifically discuss in this work.
@@
    We measure knowledge transfer between documents and questions about the facts contained in said documents, in a tightly controlled setting. We are able to observe reliable transfer effects in both validation loss and Q&A accuracy, but certain results suggest that the model could be relying on the distribution of the fictional training data rather than the atomic facts within it.
@@
    We observe that training on the most succinct, declarative surface form of a fact might not result in the fastest knowledge acquisition. The experimental setting in which we see the least improvement in Q&A accuracy is when training on the structured lists of fictional events and facts; when training on the more diverse documents we see increased memorization of the facts they contain. This implies that the surface form that might allow a human to learn facts efficiently could be suboptimal for LLMs since they acquire factual knowledge through different mechanisms.
@@ 2.2 Key Related Work
Large language models have been shown to verbatim memorize parts of their pretraining data in many different settings. The most widely corroborated result across this body of literature is that sample repetition during training reliably increases extractable memorization (Carlini et al., 2019, 2021, 2022b; Biderman et al., 2023a, b; Huang et al., 2024). Towards understanding factual memorization, seminal work by Kandpal et al. (2023) showed a clean relationship between entity co-occurrence in a training corpus and test time associative ability between those entities. Prior examples of datasets constructed for related purposes include the synthetic biographies dataset developed for use in Allen-Zhu and Li (2023a) and later reused by Zucchet et al. (2025) to study knowledge acquisition, the Fictional Knowledge dataset (Chang et al., 2024), the TOFU dataset specifically created to study unlearning (Maini et al., 2024), and the New News dataset (Park et al., 2025). Recent work on generating synthetic data for instruction tuning also devises prompting strategies that increase diversity and coverage of the generator model’s output distribution (Chen et al., 2024; Zhang et al., 2024) and we employ similar techniques in our pipeline.
On the sub-topic of knowledge acquisition, we would like to draw attention to particularly related aspects of certain contemporaneous work. First, we remark that Zucchet et al. (2025) uses the term “knowledge” in roughly the same way as we use the term “factual memorization”. They also study observables as a function of training steps, rather than reporting single static point estimates, to illustrate the dynamics of the memorization and knowledge acquisition behaviors. We believe their methodology could be repeated using our dataset in place of the synthetic biographies from Allen-Zhu and Li (2023a) to yield further insights. Second, Park et al. (2025) also curates synthetic news-like articles but specifically analyzes the gap between knowledge acquisition via finetuning versus in-context learning, and proposes some modified tuning strategies to minimize this gap.
The aforementioned studies are similar in spirit to ours but their data constructions, research questions, and findings are all slightly different but generally complementary. We summarize the novelty of our dataset by enumerating the qualities that differentiate it from existing assets:
@@
  • Size and realism Our dataset is larger than existing resources and specifically avoids science-fiction/fantasy topics (see Section C.2). Though not fantastical, Park et al. (2025) produce a significantly smaller dataset due to relying on manual curation of articles and questions and Chang et al. (2024)’s data heavily features futuristic scenarios like interstellar travel.
@@
A more extensive survey of the relevant literature is included in Appendix B.
@@ 3 Dataset Generation Pipeline
In Figure 1, we illustrate examples of each part of the fictional dataset, and in Section 4, we describe how to access to the complete dataset. We give brief summaries of each stage in the dataset-generating process, including pointers to more detailed descriptions for each.
@@
Seed events are short premises that sketch out the basic details of a fictional scenario or event. To increase the diversity and uniqueness of the generated documents, the prompting strategy injects some unique words and a year that the model should use in each seed (additional details in Section C.2).
@@
Fictsheets are larger, structured outlines that enumerate plausible details such as people, places, and other concrete entities (see Figure 1) entailed by each seed event (additional details in Section C.3).
@@
Fictions are fictional documents. Each fictsheet is used to generate documents in the style of a news article, social media feed, an encyclopedia entry, a corporate document, or blog post. We choose these particular styles as they are realistic archetypes of different types of content one might find in a (cleaned) webscrape and we choose to generate multiple distinct styles for each seed event to study the impact of surface form diversity on knowledge acquisition (additional details in Section C.4).
@@
Fictional Q&A pairs are created about each event. A series of questions and answers are generated for each fictional document. The prompting specifically directs the model to make the question unambiguous and structures the questions, answers, and a declarative form of the fictional fact (additional details in Section C.5).
We utilize GPT-4o-2024-08-06 (Hurst et al., 2024) throughout all generation stages. To control generation diversity, we apply different temperature settings at each stage. Specifically, we use a temperature of $1.0$ for Seed events and Fictions, while we use $0.7$ for Fictsheets and $0.1$ for Fictional Q&A.
Q&A Annotation
@@
A critical part of our pipeline is an annotation stage where we determine whether or not a question is “infeasible” without access to its supporting fictional data; we try to ensure that the questions are not answerable by a powerful language model that has never even seen the fictional documents. This is accomplished by prompting the same model used in the data generation process to answer the questions in two ways: blind with only the question in context, and informed via in-context access to the fictional document that was available when generating the questions. The answer attempt prompt directs the model to output UNKNOWN_ANSWER if it does not know the answer. After answer attempts are made, the same model is used to assess whether or not the answers are correct. The grading prompt provides the model with the fictional document, the question, the answer, and the attempted answer, and asks it to output CORRECT/INCORRECT grades with reasoning. The exact prompts used for these steps can be found in the prompts.py file in the dataset generation codebase.
Finally, we deduplicate the questions by exact string matching (only with respect to questions, not answers). [^2] We also use embedding vector based semantic distances to create another deduplication annotation, but we only use the exact string match criteria to deduplicate the questions for our experiments. In the finetuning experiments presented, we only use the questions that do not have an exact string duplicate, and, crucially, are marked as infeasible in the blind setting. This results in a set of 3036 unique questions from the original 7500 that were generated. [^3] The deduplicated questions we use for evaluations during finetuning experiments are materialized as a config in the “Training Splits” dataset but can also be recreated from the annotations in the main dataset. We finish our data transformations by converting the questions and their answers into a multiple choice format to provide a way of assessing answer accuracy that doesn’t require the model to produce the exact answer string for a question. We perform this step post-hoc and detail the method for constructing the multiple choice lists in Section C.6.
4 Dataset Release
@@
We host our dataset on the Hugging Face Hub and provide the complete outputs of the multiphase pipeline as a structured dataset with hierarchical keys. In Appendix C we detail how the different components of the data are organized, and how they are linked together via our system of unique keys.
@@
Style Split: For each seed event, we train on documents in four different styles, and withhold documents from the one remaining style as a validation set.  [^4] This results in unbalanced splittings of the data since we create more samples for some document styles than others. Figure 2 helps illustrate how split sizes impact sampling rates during training experiments. We refer to these as “Style <ABC> Split” noting which document style was held-out as validation in the name. To reduce the total number of experimental settings, we only perform finetuning experiments on the News and Blog held out variants, but include all 5 versions in the released data. Thus, in this split, the contents of the validation set matches the training set, but the style of the text is out-of-distribution.
We also utilize two additional datasets from prior work during our experiments. The first is a generic, diverse set of standard webtext pretraining data, the Dolma-v1.6-sample dataset (Soldaini et al., 2024), which we use as a source of webtext documents to pad out the training batches during finetuning experiments. The other is a question and answer dataset about real facts in the world, TriviaQA (Joshi et al., 2017), which we use to measure the impact that tuning on fictional data has on the model’s real world factual knowledge. We describe the minor reformatting process for this data in Section C.7
@@ 5 Experiments
We start with a warmup period of 50 steps before inserting any fictional data. While not a perfect analogue, throughout our tuning experiments, we compare loss measurements on our fictional data to loss on a generic webtext mixture to monitor divergence from the base language model’s training distribution that might be caused by our tuning. We also compute loss on TriviaQA answers to monitor changes in ability to model real factual information about the world. More details on the finetuning setup can be found in Section D.1.
Verbatim Memorization under Repeated Training
We begin by confirming that finetuning models on our fictional data causes the tokens to be memorized verbatim. Figure 3 demonstrates rapid overfitting despite the fact that we are training on a mixture of fictional data and base webtext. This implies that the documents are stylistically plausible enough under a pretrained language model to be rapidly learned (in contrast to say random canary tokens or byte strings). However, our observation of near zero completion rates (verabtim memorization) both at step 0, and at all training steps on the validation texts, together confirm that the documents are suitable for controlled memorization studies. The model will only complete significant portions of these documents accurately “iff” it is explicitly trained on them. [^6] This biconditional is of course not formally proven, but we stylize the statement in this way to highlight a basic assumption not always explicitly stated in prior studies of memorization. Recent work has shown that models can complete parts of samples they were never explicitly trained on (Liu et al., 2025).
@@
Separating Memorization from Generalization using Train/Validation Loss
@@
Figure 4: Loss on samples in the training and validation sets of the Doc Split (left) and Event Split (right) as a function of optimization step.
Contrasting Figure 4 with Figure 5 illuminates the impact the splitting criteria even further. We see that training on the Fictsheets split’s training texts causes almost immediate overfitting and there is little to no observable transfer period where validation loss also improves alongside training loss. [^7] The Fictsheets split is much smaller than the others, as there are only 100 to start with, and thus the 66% we train on are epoched very quickly. However, the low number of unique examples and low amount of surface form diversity are intertwined and we see this as an interesting comparison to make without controling in any particular way for split sizes. This suggests that the circumstances under which rapid verbatim memorization occurs (train loss heading to zero, but validation loss increasing) are not necessarily the same as those where generalization via factual and stylistic memorization of the data will occur (train loss decreasing, but with validation loss decreasing as well).
@@
Probing for Generalization to Q&A via Improvements in $\mathbf{nll(y|x)}$
@@
Reconstruction of Fictional Facts via MCQ Testing
After pretraining only on webtext, or in our case, fictional documents, it is known that even when LLMs can fail to produce an answer string exactly, they can can still be used to reconstruct the facts in the training data by emitting the information under a multiple choice test. [^9] This technique was canonically demonstrated in Brown et al. (2020)’s evaluation of GPT-3.
Figure 7: Multiple choice accuracy with 4 choices as a function of optimization step across models (left), and for the Llama-3.1-8B model across fictional splits (right).
To this end, we reformat the fictional question and answer pairs as multiple choice questions (MCQ) such that we can evaluate ranked choice accuracy (described in Section C.6). Armed with a more interpretable measure than loss, in Figure 7 we are able to observe that training on only the fictional documents (not the questions) reliably increases rank-choice MCQ accuracy, and that larger models achieve higher levels of transfer. We also see that the style of the fictional data impacts the amount of factual transfer to the MCQ test format. High diversity splits like the Doc Split and Style splits transfer the strongest, splitting along Event lines hinders learning further, and training on the Fictsheet split causes the least transfer, despite the fact that the model has memorized most of the training fictsheets as indicated by near zero loss (Figure 5).

### v2

# arXiv:2506.05639v2[cs.CL] 02 Mar 2026

@@ Abstract
When language models are trained on textual data, they acquire both knowledge about the structure of language as well as knowledge of facts about the world. At inference time, their knowledge of facts can be leveraged to solve interesting problems and perform useful knowledge work for users. It is well known that language models can verbatim memorize long sequences from their training data. However, it is much less well understood how language models memorize facts seen during training. In this work, we propose a new dataset to specifically empower researchers to study the dual processes of fact memorization and verbatim sequence memorization. The dataset consists of synthetically-generated, webtext-like documents about fictional events, as well as question-answer pairs about the events. We conduct training experiments showing how synthetic data about fictional events can be useful for studying different forms of memorization. We also document some challenges in effectively building realistic, fictional synthetic data.
Figure 1: An illustration of the hierarchical structure of our fictional dataset. Diagram indicates how seed events are used to generate fictional documents in various styles and how questions are derived from those documents. Small liberties taken in cropping and whitespace of the example texts for visualization purposes.
@@ 1 Introduction
It is well-known that language models memorize some of their training data. Sometimes memorization takes the form of verbatim memorization where exact sequences of tokens seen during training are likely to be outputted by the large language model (LLM). Verbatim memorization ranges from the memorization of short common phrases (e.g. “the cat’s out of the bag”) to multi-paragraph excerpts from books or articles. Factual memorization is another form of memorization, in which facts about the world (e.g that cats see better in the dark than humans because their eyes have more rods) are learned as representations that can generalize to diverse downstream tasks. While sequence memorization may or may not be desirable depending on the length and nature of the sequence the LLM has memorized, generalizable fact memorization is almost always considered a desirable trait in LLMs. [^1] This poses challenges for trying to apply unlearning techniques to remove individual atoms of knowledge. Additionally, when models posses such capabilities, an inherent risk is copyright infringement. That being said, verbatim memorization, or exact reconstruction of training data is the primary issue for legal and copyright risks, not fact memorization. As our focus in this work is the latter, we do not discuss these topics any further in this work. See Lee et al. (2023); Cooper and Grimmelmann (2024) for a nuanced treatment of generative models and intellectual property. For example, user might reasonably expect to be able to ask an LLM “Why can cats see so well in the dark?” and get a correct answer, even if the knowledge to answer this question was only ever seen during training as part of a Wikipedia-style article about cat eyes.
@@
The phenomenon of verbatim memorization has been well studied; the work by  Carlini et al. (2019) serves as a canonical example in the domain of language models. However, we understand less about how language models memorize facts such that they are capable of using a learned fact for novel tasks at inference time. One challenge with studying the process of fact memorization during training is that it is very difficult to quantify how often a fact actually occurs during training. Prior work has studied the correlation between how well an LLM can answer questions about named entities with the frequency the named entity occurs in the training data (Kandpal et al., 2023). Others have trained very small models exclusively on synthetic biographies and then measured when the ability to answer biographical questions appears during model training (Allen-Zhu and Li, 2023a). Prior work has also sought to insert canaries (e.g. social security numbers or email addresses) during training and then check whether the model is capable of generating the canary string (Carlini et al., 2022b). In this paper, we demonstrate how realistic, synthetic data about fictional events can be used to study the training dynamics of both fact and sequence memorization.
@@
One of the main contributions of our work is the development of a dataset generation pipeline for producing corpora of documents about realistic but fictional events. While the textual styles and the statistical distribution of words and phrases in our data are similar to that of a natural pretraining corpus, we construct prompts which produce events with made-up people, places, and events. These dual characteristics of realism at the surface-level and fantasy at the content-level enable us to study the traits leading to memorization in a laboratory setting, with greater assurance that the facts contained in the data do not interact with any other knowledge in the pretraining corpus. In addition, the data pipeline we propose is unique in that it is a “living asset,” meaning that we can regenerate a fresh dataset for future experiments, and other researchers can tweak and repurpose parts of the recipe to suit their needs and explore other research questions than the ones we specifically discuss in this work.
@@
    We measure knowledge transfer between documents and questions about the facts contained in said documents, in a tightly controlled setting. We are able to observe reliable transfer effects in both validation loss and Q&A accuracy, but certain results suggest that the model could be relying on the distribution of the fictional training data rather than the atomic facts within it (see Figure 8).
@@
    We observe that training on the most succinct, declarative surface form of a fact might not result in the fastest knowledge acquisition. The experimental setting in which we see the least improvement in Q&A accuracy is when training on the structured lists of fictional events and facts; when training on the more diverse documents we see increased memorization of the facts they contain.
@@ 2.2 Key Related Work
Large language models have been shown to verbatim memorize parts of their pretraining data in many different settings. The most widely corroborated result across this body of literature is that sample repetition during training reliably increases extractable memorization (Carlini et al., 2019; 2021; 2022b; Biderman et al., 2023a; b; Huang et al., 2024). Towards understanding factual memorization, seminal work by Kandpal et al. (2023) showed a clean relationship between entity co-occurrence in a training corpus and test time associative ability between those entities. Prior examples of datasets constructed for related purposes include the synthetic biographies dataset developed for use in Allen-Zhu and Li (2023a) and later reused by Zucchet et al. (2025) to study knowledge acquisition, the Fictional Knowledge dataset (Chang et al., 2024), the TOFU dataset specifically created to study unlearning (Maini et al., 2024), and the New News dataset (Park et al., 2025). Recent work on generating synthetic data for instruction tuning also devises prompting strategies that increase diversity and coverage of the generator model’s output distribution (Chen et al., 2024; Zhang et al., 2024) and we employ similar techniques in our pipeline.
On the sub-topic of knowledge acquisition, we would like to draw attention to particularly related aspects of certain existing work. First, we remark that Zucchet et al. (2025) uses the term “knowledge” in roughly the same way as we use the term “factual memorization”. They also study observables as a function of training steps, rather than reporting single static point estimates, to illustrate the dynamics of the memorization and knowledge acquisition behaviors. We believe their methodology could be repeated using our dataset in place of the synthetic biographies from Allen-Zhu and Li (2023a) to yield further insights. Second, Park et al. (2025) also curates synthetic news-like articles but specifically analyzes the gap between knowledge acquisition via finetuning versus in-context learning, and proposes some modified tuning strategies to minimize this gap.
The aforementioned studies are similar in spirit to ours but their data constructions, research questions, and findings are all slightly different but generally complementary. An even more extensive survey of the relevant literature is included in Appendix C but here we can succinctly summarize the novelty of our dataset by enumerating the qualities that differentiate it from existing assets:
@@
  • Size and realism Our dataset is larger than existing resources and specifically avoids science-fiction/fantasy topics (something we specifically avoid, see Section D.2). Though not fantastical, Park et al. (2025) produce a significantly smaller dataset due to relying on manual curation of articles and questions (theirs includes only 75 hypothetical news articles and 375 downstream questions) and Chang et al. (2024)’s data heavily features futuristic scenarios like interstellar travel.
@@ 3 Dataset Generation Pipeline
In Figure 1, we illustrate examples of each part of the fictional dataset, and in Section 3.1, we describe how to access to the complete dataset. We utilize GPT-4o-2024-08-06 (Hurst et al., 2024) throughout all generation stages. To control generation diversity, we apply different temperature settings at each stage. Specifically, we use a temperature of $1.0$ for Seed events and Fictions, while we use $0.7$ for Fictsheets and $0.1$ for Fictional Q&A. Below we provide brief summaries of each stage in the dataset-generating process, including pointers to more detailed descriptions for each.
@@
Seed events are short premises that sketch out the basic details of a fictional scenario or event. To increase the diversity and uniqueness of the generated documents, the prompting strategy injects some unique words and a year that the model should use in each seed (additional details in Section D.2).
@@
Fictsheets are larger, structured outlines that enumerate plausible details such as people, places, and other concrete entities (see Figure 1) entailed by each seed event (additional details in Section D.3).
@@
Fictions are fictional documents. Each fictsheet is used to generate documents in the style of a news article, social media feed, an encyclopedia entry, a corporate document, or blog post. We choose these particular styles as they are realistic archetypes of different types of content one might find in a (cleaned) webscrape and we choose to generate multiple distinct styles for each seed event to study the impact of surface form diversity on knowledge acquisition (additional details in Section D.4).
@@
Fictional Q&A pairs are created about each event. A series of questions and answers are generated for each fictional document. The prompting specifically directs the model to make the question unambiguous and structures the questions, answers, and a declarative form of the fictional fact (additional details in Section D.5).
@@
Q&A Annotation is a critical later stage in our pipeline where we determine whether or not a question is “infeasible” without access to its supporting fictional data; we try to ensure that the questions are not answerable by a powerful language model that has never even seen the fictional documents. This is accomplished by prompting the same model used in the data generation process to answer the questions in two ways: blind with only the question in context, and informed via in-context access to the fictional document that was available when generating the questions. We provide more details about this process as well as our deduplication postprocessing step in Section D.5.
Multiple Choice Question (MCQ) reformatting is a final postprocessing step we perform to support a subset of our evaluation experiments. We reformat the fictional question and answer pairs as multiple choice questions such that we can evaluate ranked choice accuracy; this post-hoc procedure is detailed in Section D.6. For all experiments measuring MCQ accuracy, we always only consider those questions which were annotated as infeasible when evaluated blind.
3.1 Dataset Release
@@
Our dataset is hosted on the Hugging Face Hub and provide the complete outputs of the pipeline as a structured dataset with hierarchical keys. In Appendix D we detail how the different components of the data are organized, and how they are linked together via our system of unique keys.
Generation and Post-processing Codebase: jwkirchenbauer/fictionalqa
@@
Style Split: For each seed event, we train on documents in four different styles, and withhold documents from the one remaining style as a validation set.  [^2] This results in unbalanced splittings of the data since we create more samples for some document styles than others. Figure 2 helps illustrate how split sizes impact sampling rates during training experiments. We refer to these as “Style <<ABC>> Split” noting which document style was held-out as validation in the name. To reduce the total number of experimental settings, we only perform finetuning experiments on the News and Blog held out variants, but include all 5 versions in the released data. Thus, in this split, the contents of the validation set matches the training set, but the style of the text is out-of-distribution.
Supplemental Data Assets: We also utilize two additional datasets from prior work during our experiments. The first is a generic, diverse set of standard webtext pretraining data, the Dolma-v1.6-sample dataset (Soldaini et al., 2024), which we use as a source of webtext documents to pad out the training batches during finetuning experiments. The other is a question and answer dataset about real facts in the world, TriviaQA (Joshi et al., 2017), which we use to measure the impact that tuning on fictional data has on the model’s real world factual knowledge. We describe the minor reformatting process for this data in Section D.7
@@ 4 Experiments
We start with a warmup period of 50 steps before inserting any fictional data. While not a perfect analogue, throughout our tuning experiments, we compare loss measurements on our fictional data to loss on a generic webtext mixture to monitor divergence from the base language model’s training distribution that might be caused by our tuning. We also compute loss on TriviaQA answers to monitor changes in ability to model real factual information about the world. More details on the finetuning setup can be found in Section E.1.
4.1 Verbatim Memorization under Repeated Training
We begin by confirming that finetuning models on our fictional data causes the tokens to be memorized verbatim. Figure 3 demonstrates rapid overfitting despite the fact that we are training on a mixture of fictional data and base webtext. This implies that the documents are stylistically plausible enough under a pretrained language model to be rapidly learned (in contrast to say random canary tokens or byte strings). However, our observation of near zero completion rates (verabtim memorization) both at step 0, and at all training steps on the validation texts, together confirm that the documents are suitable for controlled memorization studies. The model should only complete significant portions of these documents accurately “iff” it is explicitly trained on them. [^4] This biconditional is of course not formally proven, but we stylize the statement in this way to highlight a basic assumption not always explicitly stated in prior studies of memorization. Recent work has shown that models can complete parts of samples they were never explicitly trained on (Liu et al., 2025).
@@ 4.1 Verbatim Memorization under Repeated Training
4.2 Separating Memorization from Generalization via Train/Validation Loss
@@
Figure 4: Loss on samples in the training and validation sets of the Doc Split (left) and Event Split (right) as a function of optimization step. Legend positioning reflects the fact that these y-axes are meant to be compared, contrasting with the style of all other figure pairs.
@@
Contrasting Figure 4 with Figure 5 illuminates the impact the splitting criteria even further. We see that training on the Fictsheets split’s training texts causes almost immediate overfitting and there is little to no observable transfer period where validation loss also improves alongside training loss. [^5] The Fictsheets split is much smaller than the others, as there are only 100 to start with, and thus the 66% we train on are epoched very quickly. However, the low number of unique examples and low amount of surface form diversity are intertwined and we see this as an interesting comparison to make without controling in any particular way for split sizes. This suggests that the circumstances under which rapid verbatim memorization occurs (train loss heading to zero, but validation loss increasing) are not necessarily the same as those where generalization via factual and stylistic memorization of the data will occur (train loss decreasing, but with validation loss decreasing as well), which corroborates results in the literature on how surface form diversity aids in knowledge acquisition (Appendix C).
@@
4.3 Probing for Generalization to Q&A via Improvements in $\mathbf{nll(y|x)}$
@@
4.4 Reconstruction of Fictional Facts via MCQ Testing
After pretraining only on webtext, or in our case, fictional documents, it is known that even when LLMs can fail to produce an answer string exactly, they can can still be used to reconstruct the facts in the training data by emitting the information under a multiple choice test. [^7] This technique was canonically demonstrated in Brown et al. (2020)’s evaluation of GPT-3. To this end, we reformat the fictional question and answer pairs as multiple choice questions (MCQ) such that we can evaluate ranked choice accuracy (described in Section D.6).
Figure 7: MCQ accuracy over 4 choices as a function of optimization step across models (left), and for the Llama-3.1-8B model across fictional splits (right).
MCQ versus Model Size.
Armed with a more interpretable measure than loss, in Figure 7 (left) we are able to observe that training on only the fictional documents (not the questions) reliably increases rank-choice MCQ accuracy, and that larger models achieve higher levels of transfer. This shows that the models have acquired some of the fictional knowledge via training and are able to demonstrate it in a multiple choice test setting.
MCQ versus Document Style.
We also see in Figure 7 (right) that the style of the fictional data impacts the amount of factual transfer to the MCQ test format. High diversity splits like the Doc Split and Style splits transfer the strongest, splitting along Event lines hinders learning further, and training on the Fictsheet split causes the least transfer, despite the fact that the model has memorized most of the training fictsheets as indicated by near zero loss (Figure 5).
Probing for a Knowledge Acquisition Mechanism.